OpenStudy (anonymous):

The distribution of total body protein in healthy adult men is approximately Normal with mean 12.3 kg and standard deviation 0.1 kg. What is, approximately, the probability that a randomly selected healthy man has a total body protein between 12.1 and 12.2 kg?

OpenStudy (kropot72):

The z-scores for 12.1 and 12.2 are found from the following formula: \[z=\frac{X-\mu}{\sigma}\] The z-score for 12.1 is \[z _{1}=\frac{12.1-12.3}{0.1}=\frac{-0.2}{0.1}=-2.0\] Can you find the z-score for 12.2 ?